# Predictive Modeling of Power Outages during Extreme Events: Integrating Weather and Socio-Economic Factors

### v1

@@ Abstract
This paper presents a novel learning-based framework for predicting power outages caused by extreme events. The proposed approach specifically targets low-probability, high-consequence outage scenarios and leverages a comprehensive set of features derived from publicly available data sources. We integrate EAGLE-I outage records (2014–2024) with weather, socio-economic, infrastructure, and seasonal event data. Incorporating social and demographic indicators reveals underlying patterns of community vulnerability and provides a clearer understanding of outage risk during extreme conditions. Four machine learning models—Random Forest (RF), Support Vector Machine (SVM), Adaptive Boosting (AdaBoost), and Long Short-Term Memory (LSTM) —are evaluated. Experimental validation is performed on a large-scale dataset covering counties in the lower peninsula of Michigan. Among all models tested, the LSTM network achieves the lowest prediction error. Additionally, the result demonstrates that stronger economic conditions and more developed infrastructure are associated with lower outage occurrence.
@@
Power system resilience, power outage prediction, machine learning.
@@ 1 Introduction
Low-Probability High-Consequence (LPHC) events—such as floods, hurricanes, and winter storms—pose severe threats to power systems reliability and resilience. [1]. These events can cause large-scale blackouts and economic losses, as seen in Hurricane Sandy (2012), Irma (2017), and the 2021 Texas winter storm [2, 3, 4]. Several studies has explored a wide range of modeling approaches to understand and predict the relationship between extreme weather conditions and outage occurrence.
@@
For example, Eskandarpour et al. [5] predicted component damage prior to hurricanes by establishing decision boundaries using logistic regression. Spatial–temporal poisson regression has also been applied to model transmission line failure probabilities during severe weather events [6]. Although these methods offer clear interpretability, their linear assumptions limit their ability to capture the complex multi-variable interactions.
@@
ML models provide an effective framework to manage complex non-linear relationships and process substantial data volumes in real-time [7]. These models outperform traditional statistical approaches due to their ability to model nonlinearities and integrate diverse datasets, including weather, infrastructure, and socio-economic indicators. Classification-based ML studies such as [8, 9] used models like SVM and boosted classifiers to determine outage occurrence or component damage. Regression-based studies have used tree-based ensembles such as Decision Trees, RF, AdaBoost, and XGBoost to predict outage counts, durations, or affected customers [10, 11, 12]. Deep learning approaches further extend predictive capability by capturing temporal and sequential patterns in outage and weather data. Arif et al. [13] employed deep neural networks to estimate restoration times, and recurrent architectures such as recurrent neural networks (RNN) to model temporal outage progression under varying environmental conditions. These models offer strong predictive performance but require large, high-quality datasets. Some studies examined on the outages driven by routine weather or equipment failures. Jaech et al. [14] applied RNN-based models with environmental and operational data to predict outage duration in Seattle, while Sun et al. [15] used AdaBoost to estimate outage counts in Kansas. Sarwat et al. [16] showed that day-to-day weather conditions influence power outage. Our work focuses on LPHC events that generate large-scale disruptions and require fundamentally different modeling considerations.
@@
The state-of-the-art studies often encounter challenges that can compromise the accuracy of their analyses. A major limitation is the scarcity of historical data on extreme events and the associated power outages. In particular, during such events, several monitoring devices may fail, resulting in incomplete and imbalanced datasets. In [17, 18, 19], Synthetic Minority Oversampling Technique (SMOTE) has been used to address data imbalance in classification tasks in power outage studies. However, SMOTE is effective for categorical targets, it is less productive for LPHC outage prediction, which is framed as a regression problem. Another limitations arises when models are trained on a single extreme event, as they exhibit limited adaptability to concurrent incidents with complex, multi-variable interactions [20]. As shown in [21], combining data from two winter storm events instead of one improves power outage prediction. Furthermore, socio-economic features are often missing in the previous models. As demonstrated in Cruz et al. [11], socio-economic indicators such as income levels and housing characteristics provide important context for understanding community vulnerability. Counties with lower income or higher unemployment tend to have fewer resources for preparedness, which can increase outage severity and prolong restoration times. Similar observations have been reported in other studies [13, 14], where result shows that socio-economic factors contribute to prolonged outage durations and higher customer outage number. Incorporating socio-economic variables therefore enables a more comprehensive understanding of outage dynamics, supports improved prediction performance for LPHC events.
@@
In our prior work [19], we developed ML models to predict power outage probabilities, investigating the impact of socio-economic and infrastructure factors on non-LPHC outage events. This paper builds on those findings and proposes enhanced methodologies to predict power outages due to extreme weather events. The contributions of this paper are outlined as follows:
@@
1) Comprehensive Data Integration: Our model combines weather, socio-economic, and infrastructure data, along with seasonal indicators, to capture the complex factors influencing power outages. This comprehensive integration significantly enhances the predictive performance of the model.
@@
2) Innovative Use of SMOGN and KNN to Address Data Scarcity in LPHC Events: This work introduces the application of the Synthetic Minority Over-sampling Technique for Regression (SMOGN) and K-Nearest Neighbors (KNN) methods to mitigate data imbalance and missing data challenges associated with LPHC events. By adapting these techniques to the outage prediction context, the proposed framework effectively addresses data scarcity and improves the accuracy and robustness of LPHC event prediction.
@@
3) Evaluation of ML Models: We evaluate multiple ML and deep learning (DL) models, with LSTM proven particularly effective due to its ability to capture temporal patterns in time-series data. This makes our model highly suited for real-world power outage forecasting.
4) Large-Scale Validation: We validate our approach on a large-scale dataset spanning 2014 to 2022 across Michigan counties, ensuring the model’s robustness and generalizability to diverse geographic regions and socio-economic conditions.
@@ 2.1 Historical Outage Data
In this study, the EAGLE-I dataset is utilized to analyze severe power outages caused by LPHC events. The EAGLE-I dataset [22], collected and managed by Oak Ridge National Laboratory (ORNL), spans from November 2014 to December 2024. This dataset, updated every 15 minutes, aggregates data from utility companies’ public outage maps using an ETL (Extract, Transform, Load) process to estimate the number of customers affected by power outages in each county. For this analysis, we focus on the outage data relevant to Michigan. Fig. 1 illustrates the daily number of customer outages in Wayne County throughout 2024.
@@
In this context, a “customer” is defined as any entity that purchases energy from a utility under a specific tariff, rather than individual persons. Given that in the EAGLE-I, number of customers affected by power outage is recorded, and to ensure a fair comparison between counties with different sizes, it is crucial to adjust it for population differences. This can be done by collecting population data for each county from the U.S. Census Bureau [^2] https://data.census.gov/table. This approach allows us to normalize outage data relative to population size. In this way, the number of customers in the EAGLE-I dataset is divided by the population of the county scaled by 2 as shown in (1).
$P_{o}=\frac{Number\>of\>customers\>lost\>power\>*2}{Population\>of\>the\>county}$ (1)
It is estimated that, on average, each customer corresponds to approximately two residents [23].
@@ 2.2 Historical Weather Data
Historical hourly weather data was obtained from Open-Meteo, an open-source weather API that provides high-resolution meteorological information on an hourly basis  [^3] https:https://open-meteo.com/. This data source serves as the dynamic variable of the model. For instance, weather data offers insights into conditions that could lead to outages, such as wind speed and temperature. We have chosen 8 features including air temperature (F), precipitation (inch), wind speed (km/h), wind gusts (km/h), shortwave radiation ($W/m^{2}$), relative humidity (%), cloud cover (%), and surface pressure (hPa). The Fig. 2 shows the locations of weather stations across the Lower Peninsula of Michigan used in this study.
The historical weather data often has missing data, particularly during LPHC events. To address this issue, we employ K-Nearest Neighbors (KNN) imputation method describe as follows:
Algorithm 1 KNN Algorithm for Finding the Nearest Counties
@@
$k\leftarrow 5$ (Number of nearest counties to be found)
$S_{i}\leftarrow\text{Station}_{i}$ (Find the k nearest counties of $S_{i}$)
$(x_{i},y_{i})$ (Coordinates of $S_{i}$)
$D_{ij}$ (Euclidean distance between $S_{i}$ and $S_{j}$)
@@
for $k=1,\dots,j$ do
      Calculate $D_{ij}$ between $S_{i}$ and $S_{j}$ using equation (2)
      Sort distances between $S_{i}$ and all other counties in descending order
      Select the first $k$ closest counties to $S_{i}$
@@
Output: 5 nearest counties to $S_{i}$
For a given $county_{i}$, the $k$ closest counties are determined by calculating the Euclidean distance between $county_{i}$ and every other county, as described in (2).
$\text{Euclidean Distance}(S_{i},S_{j})$ $=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$ (2)
where $(x_{i},y_{i})$ and $(x_{j},y_{j})$ represent the coordinates of counties $S_{i}$ and $S_{j}$, respectively. Then, impute the missing data by averaging the selected neighbors:
$Missing\>Value=\frac{1}{j}\sum_{i=1}^{j}{Observed\>value_{i}}$ (3)
where $j$ represents the number of nearest counties considered for the imputation, identified by the K-Nearest Neighbors (KNN) algorithm [24]. Finally, the dataset is normalized using min-max scaler.
@@ 2.3 Socio-economic Data
Socio-economic indicators provide meaningful insights into community vulnerability and resilience. This hidden knowledge is instrumental in understanding factors that are not immediately obvious from basic data analysis. For this study, socio-economic indicators were obtained from the U.S. Census Bureau, including the distribution of the year when residential structures were built, the average household income, and the county-level unemployment rate. The selected data was retrieved from the U.S. Census Bureau American Community Survey (ACS) 5-Year Estimates via the Census API.  [^4] https://api.census.gov/data/2021/acs/acs5
@@ 2.4 Power Infrastructure Data
The infrastructure features, including the number of poles, towers, substations, transformers, and lines for each county, are collected from the Open Street Map dataset [25] as shown in Fig. 3. Given that the detailed topology of the power distribution network is typically not publicly available, data from the Open Street Map can serve as a useful approximation of the actual topology for research purposes. Furthermore, they provide information on how extensively the system might be affected by extreme conditions such as high winds [26].
@@
Figure 3: Power infrastructure network of counties in Lower Peninsula of Michigan.
3 Data-driven Outage Prediction Algorithm
The proposed multi-step county-level power outage prediction algorithm is illustrated in Fig. 4. The algorithm follows a four-step process. First, data preparation is performed, where different input features are preprocessed and integrated into a unified dataset. Second, the integrated dataset is rebalanced to reduce the impact of data imbalance and missing data. Third, multiple machine learning models are trained to capture the complex relationships between weather, socio-economic, and infrastructure factors and outage occurrence. Finally, the performance of the models is evaluated using standard metrics.
Figure 4: The framework of the proposed power outage prediction algorithm.
@@ 3.1 Identifying Extreme Weather Event Data
The outage and weather dataset spans the period from 1 November 2014 to 30 December 2024. Within this time frame, a total of 236 weather-related extreme outage events were identified from the National Centers for Environmental Information (NOAA)  [^5] https://www.ncdc.noaa.gov/stormevents/. Fig. 5 illustrates the distribution of extreme weather events across 20 Michigan counties.
@@
This distribution shows that extreme events are sparse. To address this limitation, seasonal information aligned with the timing of distinct extreme weather patterns is incorporated. For a specified extreme event day (denoted as $d_{ex}$), we identify all days within the same season and include them in the set $R$. This is defined as:
$R=\{d_{1},d_{2},\ldots,d_{j}\}\text{ where }d_{j}\in(m_{i-1},m_{i},m_{i+1})$
where, $m_{i}$ represents a month within a season, selected from the set $M=\{m_{1},m_{2},m_{3},\ldots,m_{12}\}$.
The next step involves calculating the $L2norm$ of the weather features between the extreme day ($d_{ex}$) and all other days in set $R$. This comparison helps identify days with weather patterns most similar to the extreme event day.
Assume $w_{d_{ex}}$ is the matrix of weather features for the extreme day (comprising 8 features, one for each weather features). The weather features for the $j$ days are represented by the matrix $W_{d}$ as follows:
$W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$
$W_{d}=\begin{pmatrix}w_{1d_{1}}&w_{2d_{1}}&\cdots&w_{8d_{1}}\\ w_{1d_{2}}&w_{2d_{2}}&\cdots&w_{8d_{2}}\\ \vdots&\vdots&\ddots&\vdots\\ w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ (4)
The L2normL2\>norm between the row vector $w_{d_{ex}}$ of $W_{d_{ex}}$ and each row vector $w_{d_{j}}$ of $W_{d}$ is calculated using the following formula:
$L2\>norm(w_{d_{ex}},w_{d_{j}})=\sqrt{\sum_{k=1}^{8}(w_{kd_{ex}}-w_{{k}d_{j}})^{2}}$ (5)
Here, $W_{d_{ex}}=\begin{pmatrix}w_{1d_{ex}}&w_{2d_{ex}}&\cdots&w_{8d_{ex}}\\ \end{pmatrix}$ is the row vector of extreme day’s weather features, $W_{d_{j}}=\begin{pmatrix}w_{1d_{j}}&w_{2d_{j}}&\cdots&w_{8d_{j}}\\ \end{pmatrix}$ is the j-th row vector of $W_{d}$, $w_{kd_{ex}}$ and $w_{{k}d_{j}}$ are the elements of $W_{d_{ex}}$ and $W_{d_{j}}$, respectively.
Subsequently, the calculation is repeated for each row in the matrix $W_{d}$, corresponding to each day $d_{j}$ within the set $R$. For each day, compute the L2normL2\>norm between the weather feature vector of the extreme day $W_{d_{ex}}$ and that of the day $d_{j}$, denoted as $W_{d_{j}}$. This results in a $j\times 1$ matrix where each element represents the $L2norm$ between $W_{d_{ex}}$ and $W_{d_{j}}$. The process is outlined mathematically as follows:
$L2\text{Norm}=\begin{pmatrix}L2\>norm(W_{d_{ex}},W_{d_{1}})\\ L2\>norm(W_{d_{ex}},W_{d_{2}})\\ \vdots\\ L2\>norm(W_{d_{ex}},W_{d_{j}})\\ \end{pmatrix}$ (6)
Finally, we select the weather features and outage data of days with the lowest $L2norm$ from matrix $L2Norm$, indicating the highest similarity to the extreme event day. Incorporating these seasonal references significantly enhances the model’s ability to differentiate between various outages in the dataset. Then, the dataset is rebalanced and used to train the machine learning models described in the following subsections.
Algorithm 2 SMOGN for Outage-Ratio Rebalancing
Input:
    $D\leftarrow$ Training dataset (hourly), with features X
    $T\leftarrow$ Threshold of minority and majority instances
    $k\leftarrow$ Number of nearest neighbors
    $O\%\leftarrow$ Percentage of over-sampling
    $U\%\leftarrow$ Percentage of under-sampling
Partition:
    $S_{min}\leftarrow$ $y_{i}\geq T$ (Minority instances)
    $S_{maj}\leftarrow$ $y_{i}<T$ (Majority instances)
/*Over-Sampling minority instances*/
for $S\in S_{min}$ do
  $\hat{S}_{min}\leftarrow|S|\times O\%$
  for $sample\in S$ do
   $KList\leftarrow$ $k$ nearest neighbors of sample
   $Distances\leftarrow$ distance between sample and each neighbor
   $SafeRange\leftarrow{0.5\times\text{median}(Distances)}$
   for $i=1$ to $\hat{S}_{min}$ do
     $X\leftarrow$ randomly selected neighbor from KListKList
     if $Distance(X,sample)<SafeRange$ then
      $X_{new}\leftarrow$ SMOTER interpolation
     else
      $X_{new}\leftarrow$ samplesample + Gaussian Noise
     end if
     Append $X_{new}$ to $D_{new}$
   end for
  end for
end for
/*Under-Sampling majority cases*/
for $S\in S_{maj}$ do
  $under\_sampled\_instances\leftarrow$ randomly sample $|S|\times U\%$ instances
  Append under_sampled_instancesunder\_sampled\_instances to $D_{new}$
end for
Output: Rebalanced training set $D_{new}$

### v2

@@ Abstract
This paper presents a novel learning-based framework for predicting power outages caused by extreme events. The proposed approach specifically targets low-probability, high-consequence outage scenarios and leverages a comprehensive set of features derived from publicly available data sources. We integrate EAGLE-I outage records (2014–2024) with weather, socio-economic, infrastructure, and seasonal event data. Incorporating social and demographic indicators reveals underlying patterns of community vulnerability and provides a clearer understanding of outage risk during extreme conditions. Four machine learning models—Random Forest (RF), Graph Neural Network (GNN), Adaptive Boosting (AdaBoost), and Long Short-Term Memory (LSTM) —are evaluated. Experimental validation is performed on a large-scale dataset covering counties in the lower peninsula of Michigan. Among all models tested, the LSTM network achieves higher accuracy.
@@
Power system resilience, power outage prediction, machine learning, extreme events.
@@ 1 Introduction
High-Impact Low-Probability (HILP) events—such as floods, hurricanes, and winter storms—pose severe threats to power systems reliability and resilience. xu2023review . These events can cause large-scale blackouts and economic losses, as seen in Hurricane Sandy (2012), Irma (2017), and the 2021 Texas winter storm buzzfeed2022 , texastribune2021 , daeli2022power . Several studies have explored a wide range of modeling approaches to understand and predict the relationship between extreme weather conditions and outage occurrence.
@@
For example, Eskandarpour et al. R13 predicted component damage prior to hurricanes by establishing decision boundaries using logistic regression. Spatial–temporal poisson regression has also been applied to model transmission line failure probabilities during severe weather events yang2019 . Although these methods offer clear interpretability, their linear assumptions limit their ability to capture the complex multi-variable interactions.
@@
ML models provide an effective framework to manage complex non-linear relationships and process substantial data volumes in real-time N30 . These models outperform traditional statistical approaches due to their ability to model nonlinearities and integrate diverse datasets, including weather, infrastructure, and socio-economic indicators. Classification-based ML studies such as N3 , eskandarpour2017leveraging used models like SVM and boosted classifiers to determine outage occurrence or component damage. Regression-based studies have used tree-based ensembles such as Decision Trees, RF, AdaBoost, and XGBoost to predict outage counts, durations, or affected customers N1 , cruz , N41 . Deep learning approaches further extend predictive capability by capturing temporal and sequential patterns in outage and weather data. Arif et al. N4 employed deep neural networks to estimate restoration times, and recurrent architectures such as recurrent neural networks (RNN) to model temporal outage progression under varying environmental conditions. These models offer strong predictive performance but require large, high-quality datasets. Some studies examined on the outages driven by routine weather or equipment failures. Jaech et al. R14 applied RNN-based models with environmental and operational data to predict outage duration in Seattle, while Sun et al. R12 used AdaBoost to estimate outage counts in Kansas. Sarwat et al. sarwat2016 showed that day-to-day weather conditions influence power outage. Our work focuses on HILP events that generate large-scale disruptions and require fundamentally different modeling considerations.
@@
The state-of-the-art studies often encounter challenges that can compromise the accuracy of their analyses. A major limitation is the scarcity of historical data on extreme events and the associated power outages. In particular, during such events, several monitoring devices may fail, resulting in incomplete and imbalanced datasets. In N26 , N14 , N17 , Synthetic Minority Oversampling Technique (SMOTE) has been used to address data imbalance in classification tasks in power outage studies. However, SMOTE is effective for categorical targets, it is less productive for HILP outage prediction, which is framed as a regression problem. Another limitations arises when models are trained on a single extreme event, as they exhibit limited adaptability to concurrent incidents with complex, multi-variable interactions N15 . As shown in N16 , combining data from two winter storm events instead of one improves power outage prediction. Furthermore, socio-economic features are often missing in the previous models. As demonstrated in Cruz et al. cruz , socio-economic indicators such as income levels and housing characteristics provide important context for understanding community vulnerability. Counties with lower income or higher unemployment tend to have fewer resources for preparedness, which can increase outage severity and prolong restoration times. Similar observations have been reported in other studies N4 , R14 , where result shows that socio-economic factors contribute to prolonged outage durations and higher customer outage number. Incorporating socio-economic variables therefore enables a more comprehensive understanding of outage dynamics, supports improved prediction performance for HILP events.
@@
In our prior work N17 , we developed ML models to predict power outage probabilities, investigating the impact of socio-economic and infrastructure factors on non-HILP outage events. This paper builds on those findings and proposes enhanced methodologies to predict power outages due to extreme weather events. The contributions of this paper are outlined as follows:
@@
1) This model combines dynamic data such as weather, seasonal indicators, historical power outages with static data such as socio-economic, and infrastructure data to capture the complex factors influencing power outages. This comprehensive integration significantly enhances the predictive performance of the model.
@@
2) This work introduces the application of the Synthetic Minority Over-sampling Technique for Regression (SMOGN) and K-Nearest Neighbors (KNN) methods to mitigate data imbalance and missing data challenges associated with HILP events. By adapting these techniques to the outage prediction context, the proposed framework effectively addresses data scarcity and improves the accuracy and robustness of HILP event prediction.
@@
3) We evaluate and compare multiple ML and deep learning (DL) models for extreme event outage prediction using integrated static and dynamic data. This makes our model highly suited for real-world power outage forecasting.
4) We train our approach on a large-scale power outage dataset across several Michigan counties. The dataset includes historical extreme events from 2014 to 2024, including the flooding event in Wayne county in June, 2020. This model can be generalized to diverse geographic regions and socio-economic conditions.
@@ 2.1 Historical Outage Data
In this study, the EAGLE-I dataset is utilized to analyze severe power outages caused by extreme weather events. The EAGLE-I dataset N25 , collected and managed by Oak Ridge National Laboratory (ORNL) on behalf of DOE, spans from November 2014 to December 2024. This dataset, updated every 15 minutes, aggregates data from utility companies’ public outage maps using an ETL (Extract, Transform, Load) process to estimate the number of customers affected by power outages in each county. For this analysis, we focus on the outage data relevant to Michigan. Fig. 1 illustrates the daily number of customer outages in Wayne County throughout 2024. In this context, a ’customer’ is defined as any entity that purchases energy from a utility under a specific tariff, rather than individual persons.
@@ 2.2 Historical Weather Data
Historical hourly weather data was obtained from Open-Meteo, an open-source weather API that provides high-resolution meteorological information on an hourly basis  [^3] https:https://open-meteo.com/. This data source serves as the dynamic variable of the model. For instance, weather data offers insights into conditions that could lead to outages, such as wind speed and temperature. We have chosen 8 features including air temperature (F), precipitation (inch), wind speed (km/h), wind gusts (km/h), shortwave radiation ($W/m^{2}$), relative humidity (%), cloud cover (%), and surface pressure (hPa). The historical weather data often has missing data, particularly during HILP events. To address this issue, KNN imputation method has been applied biswas5151316data . For a given county, the $k$ closest counties are determined by calculating the Euclidean distance between county. Then, impute the missing data by averaging the selected nearest neighbors, identified by the KNN algorithm as follows in Algorithm 1.
Algorithm 1 Algorithm to Estimate Missing Data
@@
Set of counties $\mathcal{C}=\{C_{1},C_{2},\dots,C_{N}\}$
Number of nearest counties $k$ (set to $k=5$)
Geographic coordinates $(x_{i},y_{i})$ of county $C_{i}$
$D_{ij}$ (Euclidean distance between $C_{i}$ and $C_{j}$)
@@
for each county $C_{i}\in\mathcal{C}$ do
  Initialize distance list $\mathcal{D}_{i}\leftarrow\emptyset$
  for each county $C_{j}\in\mathcal{C}(i\not=j)$ do
   Compute distance:
$D_{ij}=\sqrt{(x_{i}-x_{j})^{2}+(y_{i}-y_{j})^{2}}$
   Store $(C_{j},D_{ij})$ to $\mathcal{D}_{i}$
@@
  Sort $\mathcal{D}_{i}$ in ascending order of $D_{ij}$
  Select the first $k$ counties from $\mathcal{D}_{i}$
  Compute the average value of the $k$ neighbors
end for
Output: Impute the missing data
@@ 2.3 Socio-economic Data
Socio-economic indicators provide meaningful insights into community vulnerability and resilience. This hidden knowledge is instrumental in understanding factors that are not immediately obvious from basic data analysis. For this study, socio-economic indicators were obtained from the U.S. Census Bureau, including the distribution of the year when residential structures were built (2021 estimate), the average household income (2021 estimate), and the county-level unemployment rate (2021 estimate). The selected data was retrieved from the U.S. Census Bureau’s 2021 American Community Survey (ACS) 5-Year Estimates via the Census API.  [^4] https://api.census.gov/data/2021/acs/acs5
@@ 2.4 Power Infrastructure Data
The infrastructure features, including the number of poles, towers, substations, transformers, and lines for each county, are collected from the Open Street Map dataset N19 as shown in Fig. 3. Power infrastructure data provides valuable insights into the characteristics of each county. Given that the detailed topology of the power distribution network is typically not publicly available, data from the Open Street Map can serve as a useful approximation of the actual topology for research purposes. Furthermore, they provide information on how extensively the system might be affected by extreme conditions such as high winds N6 .
@@
Figure 3: Power infrastructure network of 20 counties in Lower Michigan.
3 Methodology
The proposed multi-step county-level power outage prediction algorithm due to extreme weather events is illustrated in Fig. 4. The algorithm follows a four-step process. First, data preparation is performed, where different input features are preprocessed and integrated into a unified dataset. Second, the integrated dataset is rebalanced to reduce the impact of data imbalance and missing data. Third, multiple machine learning models are trained on the data to capture the complex relationships between weather, socio-economic, and infrastructure factors and outage occurrence. Finally, the performance of the models is evaluated using standard metrics. In the following, we discuss the last three steps, and then the model formulation is presented.
Figure 4: The proposed learning-based framework for power outage prediction.
@@ 3.1 Identifying Extreme Weather Event Data
The outage and weather dataset spans the period from 1 November 2014 to 30 December 2024. Within this time frame, a total of 236 weather-related extreme outage events were identified from the National Centers for Environmental Information (NOAA)  [^5] https://www.ncdc.noaa.gov/stormevents/. Fig. 5 illustrates the distribution of extreme weather events across 20 Michigan counties, showing that Wayne and Hillsdale counties experienced the highest number of events.
@@
This distribution shows that extreme events are sparse. To address this limitation, seasonal information aligned with the timing of distinct extreme weather patterns is incorporated. Since HILP events are characterized by large outage magnitudes, customer outage counts are used directly as the severity indicator. Two complementary indicators are defined to identify HILP outage conditions. Let $y(c,t)$ denote the number of customers without power in county $c$ at time $t$, as reported by the EAGLE-I dataset. Since HILP events are characterized by large outage magnitudes, customer outage counts are used directly as the severity indicator.
3.1.1 Identification of HILP Outage Events
Two complementary indicators are defined to identify HILP outage conditions.
First, a weather-based extreme indicator is defined as
[…]
Second, an extreme outage indicator is defined using the empirical distribution of customer outage counts during extreme weather periods. Let $Q_{\alpha}$ denote the $\alpha$-quantile of $y(c,t)$. An outage is considered high-consequence if
$I^{\mathrm{HC}}_{c,t}=\begin{cases}1,&y(c,t)\geq Q_{\alpha},\\ 0,&\text{otherwise},\end{cases}$ (2)
where $\alpha=0.7$ is used in this study, corresponding to the top 30% of outage magnitudes observed during extreme weather conditions.
The set of HILP seed events is therefore defined as
$\mathcal{E}_{0}=\left\{(c,t)\mid I^{\mathrm{WX}}_{c,t}\cdot I^{\mathrm{HC}}_{c,t}=1\right\}.$ (3)
These seed events represent extreme outage and serve as reference points for identifying meteorologically similar extreme conditions.
3.1.2 Seasonal Analog Selection Using Weather Similarity
For each HILP seed event $(c,t_{\mathrm{ex}})\in\mathcal{E}_{0}$, a set of candidate analog periods is selected to preserve seasonal consistency. Let $m(t)$ denote the calendar month associated with time $t$. The seasonal candidate set is defined as
$\mathcal{R}(c,t_{\mathrm{ex}})=\left\{(c,t)\;\middle|\;m(t)\in\left\{m(t_{\mathrm{ex}})-1,\;m(t_{\mathrm{ex}}),\;m(t_{\mathrm{ex}})+1\right\}\right\}.$ (4)
For each county and time, an aggregated weather feature vector is constructed as
$\mathbf{x}_{c,t}=\big[T_{c,t},\,\mathrm{Prec}_{c,t},\,WS_{c,t},\,WG_{c,t},\,SWR_{c,t},\,RH_{c,t},\,CC_{c,t},\,SP_{c,t}\big]^{\top},$ (5)
where the variables represent daily or hourly aggregates of temperature, precipitation, wind speed, wind gust, shortwave radiation, relative humidity, cloud cover, and surface pressure.
To ensure balanced comparison across heterogeneous weather variables, feature-wise standardization is applied:
$\tilde{\mathbf{x}}_{c,t}=\frac{\mathbf{x}_{c,t}-{\mu}}{{\sigma}},$ (6)
where $\mu$} and $\sigma$} denote the mean and standard deviation vectors computed over the full weather dataset.
Weather similarity between a seed event $t_{\mathrm{ex}}$ and a candidate time $t$ is quantified using the Euclidean distance in standardized feature space:
$\delta(c,t_{\mathrm{ex}},t)=\left\|\tilde{\mathbf{x}}_{c,t_{\mathrm{ex}}}-\tilde{\mathbf{x}}_{c,t}\right\|_{2}.$ (7)
For each seed event, the $K$ most similar analog periods are selected as
$\mathcal{A}(c,t_{\mathrm{ex}})=\operatorname*{arg\,min}_{\begin{subarray}{c}t\in\mathcal{R}(c,t_{\mathrm{ex}})\\ |\mathcal{A}|=K\end{subarray}}\delta(c,t_{\mathrm{ex}},t).$ (8)
The final extreme-event dataset used for training is constructed as
$\mathcal{E}=\mathcal{E}_{0}\;\cup\;\bigcup_{(c,t_{\mathrm{ex}})\in\mathcal{E}_{0}}\mathcal{A}(c,t_{\mathrm{ex}}).$ (9)
This procedure expands the HILP conditions by incorporating meteorologically similar extreme-weather periods while preserving seasonal characteristics.
Algorithm 2 Data Rebalancing for High-Impact Customer Outages
1:Input:
2:    $\mathcal{T}\leftarrow$ Hourly training set with feature matrix $\mathbf{Z}$ and target $q$
3:    $\tau\leftarrow$ Outage-impact threshold
4:    $K\leftarrow$ Number of nearest neighbors
5:    $\rho_{o}\leftarrow$ Over-sampling rate
6:    $\rho_{u}\leftarrow$ Under-sampling rate
7:Partition:
8:    $\mathcal{C}_{\text{hi}}\leftarrow\{(\mathbf{z}_{j},q_{j})\in\mathcal{T}\mid q_{j}\geq\tau\}$ (High-impact cases)
9:    $\mathcal{C}_{\text{lo}}\leftarrow\{(\mathbf{z}_{j},q_{j})\in\mathcal{T}\mid q_{j}<\tau\}$ (Low-impact cases)
10:    $\mathcal{T}_{\text{bal}}\leftarrow\emptyset$
11:/* Synthetic generation for high-impact outages */
12:for $(\mathbf{z}_{r},q_{r})\in\mathcal{C}_{\text{hi}}$ do
13:  $N_{\text{syn}}\leftarrow\rho_{o}$
14:  $\mathcal{N}_{r}\leftarrow$ $K$ nearest neighbors of $(\mathbf{z}_{r},q_{r})$ in $\mathcal{C}_{\text{hi}}$
15:  $\Delta_{r}\leftarrow$ pairwise distances between $(\mathbf{z}_{r},q_{r})$ and $\mathcal{N}_{r}$
16:  $\lambda_{r}\leftarrow 0.5\times\text{median}(\Delta_{r})$
17:  for $m=1$ to $N_{\text{syn}}$ do
18:   $(\mathbf{z}_{m},q_{m})\leftarrow$ randomly selected neighbor from $\mathcal{N}_{r}$
19:   if $\mathrm{dist}\big((\mathbf{z}_{m},q_{m}),(\mathbf{z}_{r},q_{r})\big)<\lambda_{r}$ then
20:     $(\mathbf{z}^{*},q^{*})\leftarrow$ SMOTER interpolation between $(\mathbf{z}_{r},q_{r})$ and $(\mathbf{z}_{m},q_{m})$
21:   else
22:     $\mathbf{z}^{*}\leftarrow\mathbf{z}_{r}+\boldsymbol{\eta}$ (Gaussian perturbation)
23:     $q^{*}\leftarrow q_{r}$
24:   end if
25:   Append $(\mathbf{z}^{*},q^{*})$ to $\mathcal{T}_{\text{bal}}$
26:  end for
27:end for
28:/* Down-sampling low-impact outages */
[…]
30:Store $\mathcal{C}_{\text{lo}}^{\downarrow}$ to $\mathcal{T}_{\text{bal}}$
31:Output: Balanced training dataset $\mathcal{T}_{\text{bal}}$